OpenStudy (anonymous):

What's the equation of the line passing through the point (-3, 1)and parallel to y = -x - 4.

OpenStudy (aroub):

Parallel=Same slope y=mx+b , m=-1 , (-3,1) 1=-1(-3)+b 1=3+b -2=b => y=-x-2

hero (hero):

aroub XD

OpenStudy (anonymous):

y=m(x-x1)+y1 soo if it is parallel then it would have the same slope therefore it is y=-1(x+3)-4